OpenStudy (anonymous):

A square floor has an area of 480 square feet. Kristin wants to put molding around the entire perimeter of the floor. Approximately how much molding does she need? A. 720 feet B. 120 feet C. 90 feet D. 88 feet

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

well i get 88 for each side then i get 352 witch isn't and answer

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

I think it's B But I am not sure

OpenStudy (anonymous):

your looking for the perimeter of a square which is P=4s area of a square is A=s^2 you are given: A= 480 find the side: \(s= \sqrt{A}\\s=\sqrt{480}=4\sqrt{30}\) solving for \(P= 4(4\sqrt{30})=87.6 ft\) or approximately: 88 ft

OpenStudy (anonymous):

I might need your help again causei didnt pass

OpenStudy (anonymous):

i'll try, coz' my dinner is almost done and i have to go and eat it XD

OpenStudy (anonymous):

A square-shaped billboard has an area of 295 square feet. Christy wants to put lights around the entire perimeter of the billboard. Approximately how much frame does she need? A. 12 feet B. 17 feet C. 69 feet D. 74 feet

OpenStudy (anonymous):

Do the same method as I did from the previous question